Daniel has one-peso coins in the left and right pockets of his pants. If he transfers one coin from his left pocket to his right pocket, the number of coins in his right pocket would be twice of what he has on the left. However, if he transfers one coin from his right to his left, the number of coins in his pockets will be equal. How many coins does he have in his right pocket? *
Answer by ikleyn(52797)

.
Daniel has one-peso coins in the left and right pockets of his pants.
If he transfers one coin from his left pocket to his right pocket,
the number of coins in his right pocket would be twice of what he has on the left.
However, if he transfers one coin from his right to his left, the number of coins
in his pockets will be equal. How many coins does he have in his right pocket? *
~~~~~~~~~~~~~~

Let R be the number of one-peso coins in his right pocket; 

let L be the number of one-peso coins in his left pocket.


Write equations as you read the problem


    2*(L - 1) = R + 1    (1)

    L + 1 = R - 1        (2)


From the second equation, L = R-2.   Substitute it into the first equation and find R


    2*((R-2) - 1) = R + 1,

    2R - 4 - 2 = R + 1

    2R - R  = 1 + 4 + 2

       R    =   7


ANSWER.  There are 7 coins in the right pocket.

Solved.


===============


Hello, may I ask you, what the symbol * means in your post ?

I saw it in several posts; so, I think, it carries some hidden meaning . . .